OpenStudy (anonymous):

Help... please... simplify the radical expression. √(72x²)

OpenStudy (anonymous):

\[6x \sqrt{2}\] ?

OpenStudy (jim766):

\[\sqrt{72x^2}= \sqrt{36} \sqrt{2}\sqrt{x^2}\]

OpenStudy (jim766):

do you see how @Jello_Submarine got that?

OpenStudy (anonymous):

\[\sqrt{(72x ^{2})} = \sqrt{(2\times4\times9\times x ^{2}} = 2\times3\times x \times \sqrt{2} = 6x \sqrt{2}\] For clarification.

OpenStudy (anonymous):

my computer is messing up. i cant really understant much of the equation you guys are posting. How did you get the answer?

OpenStudy (anonymous):

You make 72x^2 into multipliers like 72x^2 = 2*4*9*x^2 and you notice, that you can bring out 4, 9 and x^2 from the square root and put them in front of the root, so you get 2*3*x*sqrt(2) as you can't bring out 2 from there.

OpenStudy (anonymous):

you took 4 and 9 out. how did you get 3?

OpenStudy (jim766):

sq rt of 4 = 2 sq rt of 9 = 3 2 * 3 = 6

OpenStudy (anonymous):

o i get it now! the equations finally showed up

OpenStudy (jim766):

\[\sqrt{72} = \sqrt{9}\sqrt{4}\sqrt{2}\] \[\sqrt{72}= 3 * 2 \sqrt{2}\] \[\sqrt{72}= 6 \sqrt{2}\]

OpenStudy (anonymous):

thanks!